OpenStudy (anonymous):

If r^n=a , then which of the following are true statements? Check all that apply. n1/r = a n√a = r a1/n = r ar = n

OpenStudy (anonymous):

I guessed n√a = r and a1/n = r ...

OpenStudy (solomonzelman):

\[r^n=a~~~~~~->~~~~~~~~~~~a=\sqrt[n]{r}\]

OpenStudy (anonymous):

what does that even mean? I'm so confused.

OpenStudy (anonymous):

\[r=a ^{\frac{ 1 }{n }}\] \[or`r=\sqrt[n]{a}\]

OpenStudy (anonymous):

so b and c?

OpenStudy (anonymous):

i don't think it is b \[\because~n \sqrt{a}=n*a ^{\frac{ 1 }{2 }}~and~ \sqrt[n]{a} =a ^{\frac{ 1 }{ n }}\]

OpenStudy (anonymous):

yeah I don't understand what you just wrote

OpenStudy (anonymous):

if it is written outside sqrt just as\[n \sqrt{a}~then~\it~is~different~\to~\sqrt[n]{a }\] \[n \sqrt{a}\neq \sqrt[n]{a}\]